OpenStudy (anonymous):

heelp please! Tammy has a clay vase shaped as a triangular prism with a base area of 42.6 square inches. It is one third filled with water. If there are 304 cubic inches of water in the vase, what is the height of the vase, rounded to the nearest tenth? Answer 6.5 inches 7.6 inches 14.5 inches 21.4 inches

OpenStudy (anonymous):

The formula is V= Bh... stilll working it out though...

OpenStudy (anonymous):

ok so we know that for a third of the height its 304 cubed So we will just plug this info to our formula but leave out the height since we dont know what a 3rd of the height is 42.6*height=304 height=304/42.6=7.136 Therfore the height for the amount of water in the vase is 7.136 Now that is a third of the height of the vase soooooo our height of the vase is 7.136*3=21.4